OpenStudy (anonymous):

Temperature Conversion The function 32 5 9() =+ CCF converts a temperature from C degrees Celsius to F degrees Fahrenheit. (a) Express the temperature in degrees Celsius C as a function of the temperature in degrees Fahrenheit F.

OpenStudy (anonymous):

can you rewrite you're equation. It's hard to follow

OpenStudy (anonymous):

I'm guessing its the standard... C = (F-32)*(5/9)

OpenStudy (anonymous):

The function F(C) = 9/5C +32 converts to a temperature ....

OpenStudy (anonymous):

How do I Express the temperature in degrees Celsius C as a function of the temperature in degrees Fahrenheit F.

OpenStudy (anonymous):

I would be solving for C in your problem statement which I accidently gave the answer to

OpenStudy (anonymous):

Got ya. Thanks. You are a life saver!

OpenStudy (anonymous):

Make sure you write C(F). Do you understand the algebra steps to do this?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

I'll show that real quick. We want to isolate the variable we want by itself on one side of the equatin. Thats our goal.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[F -32= {9 \over 5}C + 32 - 32\]\[(F -32)* {5 \over 9}= {9 \over 5}C * {5 \over 9}\]\[(F-32){5 \over 9}\]What you do to one side you have to do to the other. Get rid of the constants first by adding the opposite to both sides. To get rid of the coefficient in front of the variable you need to divide both sides by the coefficient. In this case it is the same thing as multiplying by the reciprocal

OpenStudy (anonymous):

did that make sense?

OpenStudy (anonymous):

There should be an "=C" in the final line. I'd write C(F)

OpenStudy (anonymous):

It makes perfect sense. I like the way you simplified this. Thanks a million.

OpenStudy (anonymous):

yw